# Many-body Fu-Kane-Mele index

### v1

@@ 5.2 Index of spin Hall states
5.2.1 Flux insertion for spin Hall states in the free fermion formalism
The charge operator in a region $\Lambda$ is given by $\Pi_{\Lambda}=(\chi_{\Lambda}\otimes\operatorname{\mathbbm{1}})\oplus(\chi_{\Lambda}\otimes\operatorname{\mathbbm{1}})$, where $\chi_{\Lambda}$ is the multiplication operator by the characteristic function of the set $\Lambda$. Let $l=(0,(-1,0))$ be the left horizontal axis so $h=h_{l}$ is the upper half plane. The analog of the many-body $U(1)$-transformation of the upper half-plane $h$ is now
@@ 5.2.1 Flux insertion for spin Hall states in the free fermion formalism
which we call the Kato generator for flux insertion. Namely, if ${\mathscr{U}}_{\phi}$ is the solution to
${\mathscr{U}}_{0}=\operatorname{\mathbbm{1}},\qquad{\mathrm{i}}\frac{{\mathrm{d}}{\mathscr{U}}_{\phi}}{{\mathrm{d}}\phi}={\mathscr{K}}_{\phi}{\mathscr{U}}_{\phi}.$ (5.19)
@@
${\mathscr{P}}_{\phi}:={\mathscr{U}}_{\phi}{\mathscr{P}}{\mathscr{U}}_{\phi}^{*}$ (5.20)
@@
corresponds to a state obtained from ${\mathscr{P}}$ by inserting a $\phi$ flux at the origin.
In particular, the Fermi projections obtained from ${\mathscr{P}}$ by inserting ±$\pm\pi$ flux in the half-line gauge are given by
${\mathscr{P}}_{\pm}={\mathscr{P}}_{\pm}^{\uparrow}\oplus{\mathscr{P}}_{\pm}^{\downarrow}\oplus 0.$ (5.21)
Since
${\mathscr{T}}{\mathscr{P}}_{-}{\mathscr{T}}^{*}={\mathscr{P}}_{+},$ (5.22)
we have that
${\mathscr{K}}{\mathscr{P}}_{-}^{\downarrow}={\mathscr{P}}_{+}^{\downarrow}{\mathscr{K}},\qquad{\mathscr{K}}{\mathscr{P}}_{-}^{\uparrow}={\mathscr{P}}_{+}^{\downarrow}{\mathscr{K}},$ (5.23)
where ${\mathscr{K}}$ is the complex conjugation111note that ${\mathscr{K}}$ leaves ${\mathcal{H}}^{\uparrow}$ and ${\mathcal{H}}^{\downarrow}$ invariant, and we identify these two spaces in this equality by $|x,\uparrow\rangle\leftrightarrow|x,\downarrow\rangle$ appearing in the definition of ${\mathscr{T}}$.
It is shown in Appendix B, Proposition B.6 that ${\mathscr{P}}^{\uparrow}_{+}-{\mathscr{P}}^{\uparrow}_{-}\in{\mathcal{J}}_{1}$ is trace class and that the Chern number of ${\mathscr{P}}^{\uparrow}$ is equal to the index of the pair of projections ${\mathscr{P}}^{\uparrow}_{+}$ and ${\mathscr{P}}^{\uparrow}_{-}$:
$\mathrm{Ch}({\mathscr{P}}^{\uparrow})=\mathrm{ind}({\mathscr{P}}_{+}^{\uparrow},{\mathscr{P}}_{-}^{\uparrow}).$ (5.24)
@@
We recall that the index of a pair of projection ${\mathscr{P}}_{1},{\mathscr{P}}_{2}$ acting on the same Hilbert space and such that ${\mathscr{P}}_{1}-{\mathscr{P}}_{2}$ is compact is given by

### v2

@@ 5 Index of $\operatorname{\mathrm{AII}}$ states
5.1 Flux insertion for $\operatorname{\mathrm{AII}}$ states in the free fermion formalism
The many-body index is defined in terms of an adiabatic flux insertion which we describe for the projection ${\mathscr{P}}$ as follows. Let $l=(0,(-1,0))$ be the left horizontal axis so $h=h_{l}$ is the upper half plane, and let $\Pi_{h}$ be the orthogonal projection onto $h$, namely $\Pi_{h}\,|x\otimes e_{i}\rangle=\chi_{h}(x)|x\otimes e_{i}\rangle$ for all $x\in{\mathbb{Z}}^{2}$ and all $i=1,\cdots,2m$. The analog of the many-body $U(1)$-transformation of the upper half-plane $h$ is now
@@ 5.1 Flux insertion for $\operatorname{\mathrm{AII}}$ states in the free fermion formalism
which we call the generator for quasi-adiabatic flux insertion. Namely, if ${\mathscr{U}}^{\mathrm{qa}}_{\phi_{2},\phi_{1}}$ is the solution of
${\mathscr{U}}^{\mathrm{qa}}_{\phi_{1},\phi_{1}}=\operatorname{\mathbbm{1}},\qquad{\mathrm{i}}\partial_{\phi_{2}}{\mathscr{U}}^{\mathrm{qa}}_{\phi_{2},\phi_{1}}={\mathscr{K}}_{\phi_{2}}{\mathscr{U}}^{\mathrm{qa}}_{\phi_{2},\phi_{1}}$ (5.6)
@@
${\mathscr{Q}}^{\mathrm{qa}}_{\phi}:={\mathscr{U}}^{\mathrm{qa}}_{\phi,0}{\mathscr{P}}\,({\mathscr{U}}^{\mathrm{qa}}_{\phi,0})^{*}$ (5.7)
@@
corresponds to a state obtained from ${\mathscr{P}}$ by inserting a $\phi$ flux at the origin. We are now interested in whether the quasi-free state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ belongs to a Kramers pair or not.
5.2 Does the state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ belong to a Kramers pair?
For any $\mu\in{\mathbb{R}}$ we denote by ${\mathscr{P}}_{\phi}^{(\mu)}=\chi_{(-\infty,\mu]}({\mathscr{H}}_{\phi})$ the Fermi projections for the Hamiltonians ${\mathscr{H}}_{\phi}$ introduced in (2.17).
The crucial step which will allow us to connect the time reversal behaviour of ${\mathscr{Q}}^{\mathrm{qa}}_{\pi}$ to the Fu-Kane-Mele index is expressing the spectral flow (2.18) defining $\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})$ as the index of the pair of projections $({\mathscr{Q}}^{\mathrm{qa}}_{\pi},{\mathscr{P}}^{(\mu)}_{\pi})$.
Proposition 5.1.
Let $\mu\in\Delta$ be such that all eigenvalues crossings of $\phi\mapsto{\mathscr{H}}_{\phi}$ across $\mu$ are simple. Then the difference ${\mathscr{Q}}^{\mathrm{qa}}_{\phi}-{\mathscr{P}}^{(\mu)}_{\phi}$ is trace class and
$\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})=\operatorname{Tr}\{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}^{(\mu)}_{\pi}\}\,\mod 2.$ (5.8)
Proof.
Since all eigenvalue crossing across $\mu$ are assumed simple, there is a finite number of values $0<\phi_{1}<\cdots<\phi_{n}<\pi$ at which crossings occur, and near any one of these values there is an eigenvalue branch locally described by a function $E_{i}(\phi)$ that is smooth and such that $E_{i}(\phi_{i})=\mu$ while $\left|\partial_{\phi}E_{i}(\phi)\right|>0$. This implies in particular that we can find $\epsilon_{i},\delta_{i}>0$ such that for all $\phi\in[\phi_{i}-\epsilon_{i},\phi_{i}+\epsilon_{i}]$ the spectrum of ${\mathscr{H}}_{\phi}$ in the interval $(E_{i}(\phi_{i})-\delta_{i},E_{i}(\phi_{i})+\delta_{i})$ consists solely and entirely of the eigenvalue $E_{i}(\phi)$, see Figure 6.
Figure 6: An eigenvalue branch $E_{i}(\phi)$ of the family $\phi\mapsto{\mathscr{H}}_{\phi}$ crossing the Fermi energy $\mu$.
For any $i=1,\cdots,n$ we then have
${\mathscr{P}}^{(\mu+\sigma_{i}\delta_{i})}_{\phi_{i}+\epsilon_{i}}-{\mathscr{P}}^{(\mu)}_{\phi_{i}+\epsilon_{i}}=\sigma_{i}{\mathscr{R}}_{i}$ (5.9)
where ${\mathscr{R}}_{i}=\chi_{(\mu-\delta_{i},\mu+\delta_{i})}({\mathscr{H}}_{\phi_{i}+\epsilon_{i}})$ is a finite rank projector and $\sigma_{i}=+1$ if $\partial_{\phi}E_{i}(\phi_{i})>0$ while $\sigma_{i}=-1$ if $\partial_{\phi}E_{i}(\phi_{i})<0$. Since there are no crossings of the line $\mu+\sigma_{i}\delta_{i}$ in the interval $[\phi_{i}-\epsilon_{i},\phi_{i}+\epsilon_{i}]$ (see again Figure 6), we have by Proposition C.6 that
${\mathscr{A}}_{i}:={\mathscr{U}}^{\mathrm{qa}}_{\phi_{i}+\epsilon_{i},\phi_{i}-\epsilon_{i}}\,{\mathscr{P}}^{(\mu+\sigma_{i}\delta_{i})}_{\phi_{i}-\epsilon_{i}}\,({\mathscr{U}}^{\mathrm{qa}}_{\phi_{i}+\epsilon_{i},\phi_{i}-\epsilon_{i}})^{*}-{\mathscr{P}}^{(\mu+\sigma_{i}\delta_{i})}_{\phi_{i}+\epsilon_{i}}$ (5.10)
is trace class and has vanishing trace. Since ${\mathscr{P}}_{\phi_{i}-\epsilon_{i}}^{(\mu+\sigma_{i}\delta_{i})}={\mathscr{P}}_{\phi_{i}-\epsilon_{i}}^{(\mu)}$ we therefore find that
${\mathscr{U}}^{\mathrm{qa}}_{\phi_{i}+\epsilon_{i},\phi_{i}-\epsilon_{i}}\,{\mathscr{P}}^{(\mu)}_{\phi_{i}-\epsilon_{i}}\,({\mathscr{U}}^{\mathrm{qa}}_{\phi_{i}+\epsilon_{i},\phi_{i}-\epsilon_{i}})^{*}-{\mathscr{P}}^{(\mu)}_{\phi_{i}+\epsilon_{i}}=\sigma_{i}{\mathscr{R}}_{i}+{\mathscr{A}}_{i}.$ (5.11)
Similarly, for $i=1,\ldots,n-1$ there are no eigenvalue crossings of ${\mathscr{H}}_{\phi}$ across $\mu$ in the interval $[\phi_{i}+\epsilon_{i},\phi_{i+1}-\epsilon_{i+1}]$ so Proposition C.6 implies that
${\mathscr{B}}_{i}:={\mathscr{U}}^{\mathrm{qa}}_{\phi_{i+1}-\epsilon_{i+1},\phi_{i}+\epsilon_{i}}\,{\mathscr{P}}^{(\mu)}_{\phi_{i}+\epsilon_{i}}\,({\mathscr{U}}^{\mathrm{qa}}_{\phi_{i+1}-\epsilon_{i+1},\phi_{i}+\epsilon_{i}})^{*}-{\mathscr{P}}^{(\mu)}_{\phi_{i+1}-\epsilon_{i+1}}$ (5.12)
is trace class and has vanishing trace.
Finally, again by Proposition C.6 and the fact that there are no level crossings of ${\mathscr{H}}_{\phi}$ across $\mu$ in the intervals $[0,\phi_{1}-\epsilon_{1}]$ and $[\phi_{n}+\epsilon_{n},\pi]$, we have that
${\mathscr{B}}_{0}$ $:={\mathscr{U}}^{\mathrm{qa}}_{\phi_{1}-\epsilon_{1},0}\,{\mathscr{P}}\,({\mathscr{U}}^{\mathrm{qa}}_{\phi_{1}-\epsilon_{1},0})^{*}-{\mathscr{P}}^{(\mu)}_{\phi_{1}-\epsilon_{1}}$ (5.13)
${\mathscr{B}}_{n+1}$ $:={\mathscr{U}}^{\mathrm{qa}}_{\pi,\phi_{n}+\epsilon_{n}}\,{\mathscr{P}}^{(\mu)}_{\phi_{n}+\epsilon_{n}}\,({\mathscr{U}}^{\mathrm{qa}}_{\pi,\phi_{n}+\epsilon_{n}})^{*}-{\mathscr{P}}^{(\mu)}_{\pi}$ (5.14)
are trace class and have vanishing traces. This together with a repeated application of (5.11) and (5.12) yields
$\operatorname{Tr}\left\{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}-{\mathscr{P}}^{(\mu)}_{\pi}\right\}=\sum_{i=1}^{n}\sigma_{i}\operatorname{Tr}R_{i}=\operatorname{\mathrm{SF}}_{\mu}([0,\pi]\ni\phi\mapsto{\mathscr{H}}_{\phi})$ (5.15)
where we use cyclicity of the trace and the fact that ${\mathscr{U}}^{\mathrm{qa}}_{\phi^{\prime\prime},\phi^{\prime}}{\mathscr{U}}^{\mathrm{qa}}_{\phi^{\prime},\phi}={\mathscr{U}}^{\mathrm{qa}}_{\phi^{\prime\prime},\phi}$ for any $\phi,\phi^{\prime},\phi^{\prime\prime}\in{\mathbb{R}}$. The claim now follows from (2.18). ∎
Proposition 5.2.
The state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ is a Kramers singlet if $\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})=0$ while it belongs to a Kramers pair if $\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})=1$.
Proof.
@@ 5.2 Does the state $\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ belong to a Kramers pair?
We recall that the index of a pair of projections ${\mathscr{P}}_{1},{\mathscr{P}}_{2}$ acting on the same Hilbert space and such that ${\mathscr{P}}_{1}-{\mathscr{P}}_{2}$ is compact is given by